OpenStudy (anonymous):

What is the slope of the line below? If necessary, enter your answer as a fraction in lowest terms, using the slash ( / ) as the fraction bar. Do not enter your answer as a decimal number or an equation.

OpenStudy (anonymous):

Welcome to OpenStudy! Is there a picture to go with this?

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

Okay, could you post it? :)

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

What is the slope of the line below? If necessary, enter your answer as a fraction in lowest terms, using the slash ( / ) as the fraction bar. Do not enter your answer as a decimal number or an equation. http://media.apexlearning.com/Images/200706/30/55ec5d48-11cd-41e0-981f-0719c4731fe8.gif

OpenStudy (anonymous):

Sorry, I didn't get a notification that you replied. Do you know the slope formula?

OpenStudy (anonymous):

Y2-Y1 / X2-X1

OpenStudy (anonymous):

@user3

OpenStudy (anonymous):

Yep, so we just substitute the points given into that formula

OpenStudy (anonymous):

ohkay after you substitute the points into the formula i get lost after that @user3

OpenStudy (anonymous):

Okay, after substituting we get \(\dfrac{7 -(-3)}{3-(-2)}\). Subtracting a negative number is the same as adding, so now we have \(\dfrac{7+3}{3+2}\). Now we add the top and bottom.

OpenStudy (anonymous):

so it will be 

OpenStudy (anonymous):

but dont you simplify it @user3

OpenStudy (anonymous):

Yes, we need to simplify \(\dfrac{10}5\)

OpenStudy (anonymous):

so the answer is m=2 @user3

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

If you have any questions about the site, feel free to let me know :)

OpenStudy (anonymous):

ohkay thanks @user3

OpenStudy (anonymous):

No problem